nolegirl24:

What is the product of Seven square root eight times four square root five? Simplify your answer. Twenty eight radical ten Fifty six radical two Fifty six radical ten Two hundred eighty radical four

Vocaloid:

first you would just multiply the square root parts together

mikewwe13:

7√8*4√5 =

Vocaloid:

sqrt(A) * sqrt(B) = sqrt(A*B) sqrt(8) * sqrt(5) = ?

nolegirl24:

40?

nolegirl24:

Wait what?!?!?

Vocaloid:

good, sqrt(40) now we just multiply the non-radical parts to get 7 * 4 = 28

Vocaloid:

now we go back to our sqrt(40) and see if we can factor out any perfect squares

Vocaloid:

40 = 4*10 sqrt(40) = sqrt(4)*sqrt(10) sqrt(4) = ?

nolegirl24:

square root of 4 is 2

Vocaloid:

good, so sqrt(40) = sqrt(4)*sqrt(10) = 2*sqrt(10) so let's put everything together 28* 2*sqrt(10) = 56sqrt(10) = C